(a) A reversible heat pump is used to maintained a temperature of 0 C in a
refrigerator, when it rejects the heat to the surrounding at 15 C. if the heat removal
rate from the refrigerator is 1440 kJ/min, determine the COP of the machine and
work input required.

Grade:

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
The COP would be given by:
COP=\frac{15}{15-0}=1
Also, we know that:
COP=\frac{Q_{i}}{W}=\frac{1440 kJ/min}{W}\Rightarrow W=1440 kJ/min
Regards
Sumit